If You deposit money into an account that pays 4.5% annual interest. If the account is compounded YEARLY, how long will it take for you account to double in value? If the account is compounded CONTINUOUSLY, how long will it take for the account to double in value?
Answer by Alan3354(30993)

If You deposit money into an account that pays 4.5% annual interest. If the account is compounded YEARLY, how long will it take for you account to double in value? If the account is compounded CONTINUOUSLY, how long will it take for the account to double in value?
-----------
Yearly:
2*Amt = Amt*(1.045)^n
1.045^n = 2
n*log(1.045) = log(2)
n = log(2)/log(1.045)
n =~ 15.747 years
----------
Continuous:
2 = e^(rn) = e^0.045n
ln(2) = 0.045n
n = ln(2)/0.045
n =~ 15.403 years